OpenStudy (anonymous):

A drawer conatins 4 red socks and 5 blue socks. 3 socks are drawn. determine the probability that: (a) exactly 2 socks are selected (b) atleast 2 red socks are drawn

OpenStudy (anonymous):

a) Exactly two socks. Two socks of what color?

OpenStudy (anonymous):

any two socks

OpenStudy (anonymous):

Can someone help me with this question?

OpenStudy (anonymous):

@mathstudent55 can you please help me with this question?

OpenStudy (anonymous):

@kropot72 can you please help me with this question?

OpenStudy (anonymous):

@ryan123345 can you help me with this question?

OpenStudy (kropot72):

Given that 3 socks are drawn, the first part of the question could be asking the probability that there is a pair of the same color that can be selected from the three.

OpenStudy (kropot72):

\[P(2\ red)=\frac{4C2\times 5C1}{9C3}\] \[P(2\ blue)=\frac{5C2\times 4C1}{9C3}\] To find the probability that there is a pair of the same color that can be selected from the three add the results: P(2 red) + P(2 blue) = ?